The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 32.6 for a sample of size 20 and a standard deviation of 8.8. Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 98% confidence level). Assume the data is from a normally distributed population.

Enter your answer as a tri-linear inequality accurate to three decimal places.

< μ <
Answer #1

Solution :

Given that,

sample size = n = 20

Degrees of freedom = df = n - 1 = 20 - 1 = 19

t\alpha /2,df = 2.539

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.539 * (8.8 / \sqrt 20)

= 4.996

The 98% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

32.6 + 4.996 < \mu < 32.6 + 4.996

27.604 < \mu < 37.596